OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). (2 points) y = 1 over 3 x + 7 y = -3x + 3 y = 1 over 3 x + 3 y = -3x + 7

OpenStudy (anonymous):

Parallel lines have the same slope. The equation of a line is y=mx+b where m is the slope and b is where the line crosses the y axis. In the given equation y=-3x+12, the slope is -3. So we are looking for an answer where the slope is -3. The equation will look something like this: y=-3x+b Answers two and four have slopes that are -3. So they are possible correct answers. To find out which one is the correct answer, you substitute the given point value(-1,6) into the equation y=-3x+b and solve for b. 6=-3(-1) +b Solve for b 6=3+b 3=b So the equation with the correct answer is y=-3x +3 So answer two is correct